A horse exerts a force of 200 lb. while pulling a sled for 3 miles. If the trip takes 30 minutes, what is the power output of the horse in horsepower?

1
Expert's answer
2022-06-07T17:37:08-0400

Given:

"3\;\rm mi=15840\: ft"

"1\rm\: lb\cdot ft/min=33000\: hp"


"P=\frac{W}{t}=\frac{Fd}{t}"

"P=\rm\frac{200\: lb*15800\: ft}{30\: min}=105333\: lb\cdot ft/min=3.19\:hp"